Problem 16-1 EBIT and Leverage [LO1] Ghost, Inc., has no debt outstanding and a total market value of $250,000. Earnings before interest and taxes, EBIT, are projected to be $42,000 if economic conditions are normal. If there is strong expansion in the economy, then EBIT will be 18 percent higher. If there is a recession, then EBIT will be 30 percent lower. The company is considering a $100,000 debt issue with an interest rate of 8 percent. The proceeds will be used to repurchase shares of stock. There are currently 10,000 shares outstanding. Ignore taxes for this problem. a-1. Calculate earnings per share (EPS) under each of the three economic scenarios before any debt is issued. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) a-2. Calculate the percentage changes in EPS when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.) b-1. Calculate earnings per share (EPS) under each of the three economic scenarios assuming the company goes through with recapitalization. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) b-2. Given the recapitalization, calculate the percentage changes in EPS when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
a-1
EPS = EBIT*(1-tax rate)/shares outstanding
Recession
EPS = EBIT*(1-recession impact%)*(1-tax rate)/shares outstanding
EPS=42000*(1-0.3)*(1-0)/10000
EPS=2.94
Normal
EPS = EBIT*(1-tax rate)/shares outstanding
EPS=42000*(1-0)/10000
EPS=4.2
Expansion
EPS = EBIT*(1+Growth impact%)*(1-tax rate)/shares outstanding
EPS=42000*(1+0.18)*(1-0)/10000
EPS=4.96
a-2
%age change in EPS for Recession
=(EPS recession/EPS normal-1)*100
=(2.94/4.2-1)*100
=-30%
%age change in EPS for Growth
=(EPS Growth/EPS normal-1)*100
=(4.956/4.2-1)*100
=18%
b-1
New no. of shares = old shares-debt/(Market value/old shares)
=10000-100000/(250000/10000)
=6000
EPS = (EBIT-debt*interest%)*(1-tax rate)/new shares outstanding
Recession
EPS = (EBIT*(1-recession impact%)-debt*interest %age)*(1-tax rate)/new shares outstanding
EPS=(42000*(1-0.3)-100000*0.08)*(1-0)/6000
EPS=3.57
Normal
EPS = (EBIT-debt*interest%)*(1-tax rate)/new shares outstanding
EPS=(42000-100000*0.08)*(1-0)/6000
EPS=5.67
Expansion
EPS = (EBIT*(1+growth impact%)-debt*interest %age)*(1-tax rate)/new shares outstanding
EPS=(42000*(1+0.18)-100000*0.08)*(1-0)/6000
EPS=6.93
b-2
%age change in EPS for Recession
=(EPS recession/EPS normal-1)*100
=(3.5667/5.6667-1)*100
=-37%
%age change in EPS for Growth
=(EPS Growth/EPS normal-1)*100
=(6.9267/5.6667-1)*100
=22%